OpenStudy (anonymous):

g(x)=(x+2)^3 solve for x? and than tell me what g^-1(x)=?

OpenStudy (anonymous):

They want you to find the inverse of the function. Do you know how to do that? @rissalee122

OpenStudy (anonymous):

yes that would be (x+2)^3=y correct

OpenStudy (anonymous):

What you do is you say ok so y = (x + 2)^3, if you want to find the inverse of this function, you switch the y and x so it becomes: x = (y + 2)^3. Now you just solve for y to get the inverse.

OpenStudy (e.mccormick):

Solving for x and getting the inverse are related. What you do with something like this is say: \[y=(x+2)^3\] Then you do operations until you have \(x=\) something.

OpenStudy (anonymous):

@rissalee122 Can you do that or you do you want me to take you through? It should be pretty straight-forward from here.

OpenStudy (e.mccormick):

@genius12 If her instructions specifcally state to solve fr x first, then doing the swap first would not ger the proper answer for that part of the question. The only difference is if you swap x and y first or last, but some teachers are picky...

OpenStudy (anonymous):

First I was to solve for y then x so I think I see now thank you both

OpenStudy (e.mccormick):

Yah, there is no difference in the work done. However, if they specifically state in two questions, "Solve for x" then "Give the inverse of..." you have no choice but to swap at the end. So like you guessed, it is the same basic thing.